NPP Projection Close 10/30

To do this evening/ was completed:

Presentation Issues:

Preshow Clips/Playlist:

Trailer Updates:

Special Events/ AV set up: Ran some test film off the platter with Alec and Lillian in booth 1...cleaned the projector and the brain/rollers after the film ran out...turned the projector breakers/tree power off...I moved the test film from the middle platter to the bottom one in preparation for tomorrow's Bugonia tech (note: I brought the pucks down and placed them around the print and placed the brain on the top platter)

Night of the Living Dead screening went smoothly...the band made some final remarks before packing up and I put the rest of the gear away back beneath the curtain in theater 1 once they departed (note: some wires may be out of place...I was not sure which cables had been taken from 2 or 4/6 but these are all split between the bins in 1...JBLs are back in 2 along with their power cords and XLRs)

Ordering/Receiving (posters & trailers): Additional posters for Keeper and FNAF 2 were delivered + one sheets for Nouvelle Vague, KPop Demon Hunters, Secret Agent, and How the Grinch Stole Christmas 25th Anniversary were delivered and stored behind 3/7

Ordering/Receiving (content):

Outgoing/Shipping (content):

Print/video Inspection: Continued inspecting Jaws in booth 1...one reel left...I left the penultimate reel on the workbench since I took the liberty of winding them all back tails out and need to finish rewinding this one

Asset tracking doc updated/checked?:

Maintenance/cleaning:

Cue sheets updated/checked (every shift):

Marquee updated?:

Posters updated?: Removed Roofman from the front and replaced it with Nouvelle Vague under a "now showing" snipe + placed a "now showing" snipe above Bugonia

Cleaning lights on for the night?: Yes

Projection cart by entrance?: Yes

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will